Calculate the spin-only magnetic moment, us (a.k.a. p/pe), for the complex [Co(CN).). Number 2.83 BM
Answer #1

spin- only magnetic moment, u= In (n+2) here, n = number of unpaired electrons. (co (CN) 73 ON= x-6=-3 __ x=+3 cot3 do All el